OpenStudy (anonymous):

solve the equation log(base 6)(x+9)-log(base 36) x=1

OpenStudy (anonymous):

try rewriting your statement in forms of the natural log, using the base change law of logarithms.

OpenStudy (anonymous):

so what you are saying is changing it to log(base6)(x+9)/log(base36)x =1 ?

OpenStudy (anonymous):

well you didn't change anything right there, but you can write log_6(x+9) as: \[\log_6(x+9)=\frac{ \ln(x+9) }{\ln6 }\]

OpenStudy (anonymous):

okay. understand that. so would you do the same with the other one too. so it will look like lnx/ln38

OpenStudy (anonymous):

Is that your problem? \[\log_6(x+9)-\log_{36}x=1\] Sorry my windows machine loads slow, so I maybe have misinterpreted the brackets in your opening post.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then yes indeed, continue that way. just notice that 36 = 6*6 = 6^2

OpenStudy (raden):